# apndxAdditional References A Hitchhiker’s Guide to Privacy-Preserving Cryptocurrencies: A Survey on Anonymity, Confidentiality, and Auditability

### v1

@@ 4.3.6. Plausible Deniability
With plausible deniability, we intend whether a user can deny intentional participation in a transaction, preventing others and adversaries from proving their involvement (e.g., (Almashaqbeh and Solomon, 2022; Ziegeldorf et al., 2018)). Existing work is roughly evenly divided between those that support this property (e.g., (Noether et al., 2016; Poelstra, 2016; Narula et al., 2018; Kiayias et al., 2022; Tomescu et al., 2022; Wüst et al., 2022; Goodell et al., 2023; Papadoulis et al., 2024)) and those that do not (e.g., (Maxwell, 2013; Van Saberhagen, 2013; Garman et al., 2017; Bünz et al., 2020; Gross et al., 2021; Baudet et al., 2023; Guan et al., 2022; Yu et al., 2024)). Basically, protocols “on-demand”, such as mixing protocols, subsume an explicit user intention in using the additional service to obtain a privacy feature (such as unlinkability). Likewise, overlay systems implementing an anonymous cryptocurrency, such as Zerocoin (Miers et al., 2013), Zerocash (Ben Sasson et al., 2014), EZC (Androulaki and Karame, 2014), Zether (Bünz et al., 2020), require explicit intention to convert native currencies to a privacy-preserving one. Conversely, systems that force anonymity or unlinkability as default, readily provide the property of plausible deniability. Among these systems, we mention Monero (Noether et al., 2016), MimbleWimble (Poelstra, 2016), UTT (Tomescu et al., 2022), PEReDi (Kiayias et al., 2022), or Platypus (Wüst et al., 2022), that do not provide the functionality of combining anonymous and public transactions.
@@ 4.4. Auditability
Auditability refers to the ability to examine, track, and verify transactions or balances while preserving a certain degree of privacy. It is a key feature of the third generation of privacy-preserving cryptocurrencies and, in particular, CBDCs. The concept of auditability is closely related to that of policy, which refers to the rules enforced by the protocol or by users to govern behaviors (e.g., to restrict who can transact and under what conditions). While policies involve ex-ante enforcement (i.e., preventing undesired actions before they occur), auditability supports post-facto verification by auditors and regulators. We point out that accountability is also important, as it implies linking actions to individuals, e.g., through identity-binding mechanisms such as those provided by a registration authority to comply with AML/CFT regulations.
Full privacy often conflicts with compliance: if transactions are completely opaque, regulators have no oversight (Ballaschk and Paulick, 2021). Some privacy designs therefore incorporate policy-controlled privacy. We identify the following approaches that enable or support auditability: anonymity budget, operating limits, revocable anonymity and user tracing, coin tracing, and verification without disclosure.
@@ 4.4.1. Anonymity Budget
An anonymity budget limits the participant’s ability to perform anonymous payments. Each transaction consumes a portion of this budget, and once the participant runs out of it, he loses the ability to make further anonymous payments (e.g., (Camenisch et al., 2006; Wüst et al., 2019; Tomescu et al., 2022; ECB, 2019)). Camenisch et al. (Camenisch et al., 2006) introduce the bounded-anonymity business model, where there is a publicly-known limit to the number of coins a user may anonymously transfer to a specific merchant. PRCash (Wüst et al., 2019) restricts the total amount of anonymous payments a user can receive within a time window (or epoch). The anonymity budget is a special non-spendable output. By adjusting the budget size and duration, authorities can regulate the flow of anonymous funds. Once the budget is exhausted, or if the user opts for non-anonymous transactions, the payee creates an output that includes his identity encrypted with the regulator’s public key. UTT (Tomescu et al., 2022) employs the anonymity budget to balance privacy with accountability for spending operations. Beyond the allocated budget, transactions may either lose privacy or require auditing. An auditor periodically refreshes anonymity budgets and can approve transactions that exceed the budget, e.g., by requesting additional transaction details before signing. In all these systems, the anonymity budget is implemented as a special token that must be included in any anonymous but accountable transaction.
@@ 4.4.2. Operating Limits
Operating limits indicate constraints beyond which transactions cannot be carried out. Limits are usually imposed on the single transaction value, overall user’s balance, amount of sent or received funds, and number of performed transactions within a time window. Compliance with limits is generally proven in zero-knowledge. Garman et al. (Garman et al., 2017) extend Zerocash by introducing several building blocks to support auditability, including regulatory closure (i.e., restricting transactions to assets of the same type), spending limits, selective user inquiries, and coin tracing. A central idea is the enforcement of a per-transaction spending limit, whereby a transaction becomes invalid if it exceeds the defined limit, unless it is signed by an authorized party. Another proposal involves maintaining a per-user counter incremented with each outgoing transaction. The system can then reject transactions if the user’s counter is above a predefined threshold. Building on these ideas, Gross et al. (Gross et al., 2021) propose a two-layered account-based system for implementing a CBDC. To enforce the spending policy, each account maintains an epoch turnover, representing the total amount spent during the current epoch. Compliance with spending limits is proven using ZKPs as part of each account update transaction. Similarly, AQQUA (Papadoulis et al., 2024) uses multiple counters per account, which are publicly stored as three commitments: one to the account balance, and two to the cumulative amounts sent and received. AQQUA also introduces a registration authority, which links users’ real-world identities to their (possibly, multiple) public keys, and an audit authority, which verifies compliance with defined policies. During an audit, the authority selects a user by his public key and requests the opening of two historical account state snapshots (for every user’s accounts). These openings allow the authority to directly inspect whether the user complies with specific policies. Five types of policies exist: spending limits, receiving limits, transaction value limits, non-participation (i.e., inactivity over a time period), and transaction opening (i.e., selective disclosure of the value sent or received in a specific transaction). After the audit, the user randomizes his accounts (and provides a ZKP of valid update), thereby restoring his privacy. Xue et al. (Xue et al., 2023) use per-user counters, as well. They design an unspent token-based ledger, which stores counters in a Merkle tree managed by regulators, which are also in charge of user on-boarding. Users can define policies using ZKPs161616Specifically, zk-SNARKs are used for arithmetic statements and Sigma Protocols for algebraic statements. on counters, which are updated for each transaction through user-regulator interaction. Counters enable policies on total amount and frequency of fund transfers within a time period. When suspicious transactions are detected, the participants’ identities can be recovered by regulators. Similarly, Chen et al. (Chen et al., 2020) consider three policies—on transaction limits, rates, and value—and recur to ZKPs to attest compliance.
@@ 4.4.3. Revocable Anonymity and User Tracing
User tracing mechanisms enable authorities to discover the user’s identity or to identify all transactions associated with a specific user when needed. This concept is often related to revocable or conditional anonymity. Revocable anonymity refers to the privacy model where users remain anonymous by default, but their identity can be revealed under certain conditions or rules. Conversely, user tracing refers to technical mechanisms that allow authorized entities to identify users or link their transactions when required.
@@
Some contributions use special digital signatures that reveal the sender’s identity if he double spends (e.g., (Hayes, 1990; Zhang et al., 2019; Beer et al., 2024)). Other approaches encrypt user-related information—such as identifiers, keys, or transaction metadata—within the transaction, in special tags, making them accessible only to tracing authorities with decryption capabilities (e.g., (Cecchetti et al., 2017; Garman et al., 2017; Zhang et al., 2019; Liu et al., 2022; Lin et al., 2020; Li et al., 2021; Kiayias et al., 2022)). This is often referred to as the trapdoor approach, where decryption is enabled through a private key or other privileged information. Many systems also rely on registration authorities that link real-world identities to cryptographic keys, making traced identities resolvable (e.g., (Zhang et al., 2019; Liu et al., 2022)). For example, Cecchetti et al. (Cecchetti et al., 2017) and Liu et al. (Liu et al., 2022) consider a setting with a central bank. While Cecchetti et al. (Cecchetti et al., 2017) manage in oblivious RAM the transactions across users’ commercial banks, Liu et al. (Liu et al., 2022) require that each transaction includes the user’s identity encrypted with the central bank public key. In a more general setting, Zhang et al. (Zhang et al., 2019) introduce a supervision authority that can extract the payer’s identity from any transaction using its private key (and the registration authority). Garman et al. (Garman et al., 2017) propose a tracing mechanism where each user receives a unique encryption key: if the user is being traced, the key is a randomized version of the authority’s key; otherwise, it is derived from a null key. Users cannot tell which type of key they received. To enhance transparency, periodically the system spends tokens that publicly demonstrate how many users were traced without revealing their identities. Androulaki et al. (Androulaki et al., 2020) assign each user to an auditor at registration time; then, the system encrypts the transaction information under the public keys of the sender’s and the receivers’ auditors. Xue et al. (Xue et al., 2023) encode the payer’s identity in each transaction using the regulator’s key and the transaction’s epoch identifier. The epoch identifier enables temporal scoping, which limits the impact of key compromising and can enable granular tracing.
A few works assign long-term keys to users, from which they derive one-time keys for individual transactions. User tracing mechanisms aim to recover the long-term keys that uniquely identify a user. For example, Lin et al. (Lin et al., 2020) propose a system in which users perform transactions using anonymous keys derived from their long-term public addresses. To enable tracing, a designated manager can use his private key and the anonymous transaction key to recover the user’s long-term public address (and his identity). Traceable Monero (Li et al., 2021) introduces a method for generating one-time key pairs for payees, where each one-time address is associated with a tag. This tag can be decrypted by the tracing authority to recover the payee’s long-term public key. The tag is designed to support both backward and forward tracing: it can identify both the accounts that sent funds to the current account and the subsequent accounts that receive funds originating from it.
The ability of tracing authorities to decrypt transactions places a high trust burden on a centralized entity. To mitigate this, Yu et al. (Yu et al., 2024) require collaboration between two distinct authorities to detect correlations among a batch of transactions—specifically, those originating from the same user—and to trace user identities when necessary. To enable this functionality, the authors introduce the concept of a mark171717Although the concepts of tag and mark are closely related, we follow the terminology of (Yu et al., 2024), using mark to denote data that enables a two-party computation protocol in a privacy-preserving manner., which embeds the sender’s public key encrypted with random values. These values can be reconstructed only through the joint effort of both authorities. Conversely, PEReDi (Kiayias et al., 2022; Sarencheh et al., 2022) proposes distributing the decryption power across multiple parties, requiring them to cooperate and reach a quorum in order to access private information. This very idea is then also used by Sarencheh et al. (Sarencheh et al., 2025).
@@ 4.4.4. Coin Tracing
Garman et al. (Garman et al., 2017) propose an approach to trace certain coins, avoiding an extensive transaction auditing. In particular, the scheme enables tracing by marking specific coins as traceable, i.e., tainted coins181818A tainted coin is a traceable digital asset, which could be linked to illicit or suspicious activity, often flagged for enhanced scrutiny or restricted use.. When such a coin is spent, all resulting output coins also become traceable. To implement this, each coin contains a fresh encryption key, and tracing data for the outputs is encrypted under the input coin’s key and under the authority’s public key to prevent the sender from performing unauthorized tracing. To limit cascade tracing and collusion risk (i.e., when a user and the authority jointly trace non-tainted coins), exchanges are allowed to verifiably remove tracing using a public dummy key. Keller et al. (Keller et al., 2021) propose a method to de-anonymize users in CoinJoin and ring signature schemes. They model a scenario where a law enforcement agency collaborates with certain users—referred to as witnesses—who can confirm that their inputs were not used in a given mixing transaction. As the number of such witnesses increases, the anonymity set of the remaining users is reduced, making de-anonymization more feasible.
@@ 4.4.5. Verification without Disclosure
Auditability mechanisms that enable verification without disclosure allow a party to confirm compliance with a policy without learning any private data. These mechanisms rely on cryptographic proofs (usually NIZK) to show that hidden data satisfies specific constraints (e.g., (Narula et al., 2018; Wüst et al., 2022; Cecchetti et al., 2017; Bünz et al., 2020; Ma et al., 2021)). Crucially, the verifier learns nothing beyond the validity of the claimed statement. Besides the approaches we already mentioned in Operating limits, here we review those supporting policies beyond basic balance or transfer checks. For example, zkLedger (Narula et al., 2018) uses Pedersen commitments enabling aggregate computations. zkLedger supports complex auditing via a map-reduce model: banks generate commitments over filtered and transformed data (e.g., checking for non-zero values or squaring for variance) along with NIZKs proofs of correctness, enabling auditors to compute metrics like means and variances. Platypus (Wüst et al., 2022) defines account states with opaque auxiliary data to support custom policy checks. Solidus (Cecchetti et al., 2017) uses PVORM and ZKPs to publicly verify a bank’s transaction log compliance. Zether (Bünz et al., 2020) implements confidential transactions in Ethereum, using NIZKs to verify properties like non-negative balances. DSC (Ma et al., 2021) enforces compliance of encrypted transactions with range and authorization proofs.
@@ 4.5. Summary and Takeaways
The analysis presented in this section highlights the diverse and evolving landscape of privacy goals in digital currencies. While early systems primarily targeted anonymity and unlinkability through address obfuscation and mixing techniques, more recent designs emphasize stronger notions of confidentiality and auditability. These are supported by advanced cryptographic tools such as ZKPs, homomorphic commitments, and threshold encryption.
Each privacy goal—anonymity, confidentiality, unlinkability, and auditability—requires distinct mechanisms and introduces specific trade-offs. For instance, unlinkability may conflict with auditability, and encrypted transactions that ensure confidentiality can complicate regulatory oversight. No single protocol fully satisfies all privacy and compliance objectives; rather, systems tend to prioritize different goals based on their threat models, trust assumptions, and intended deployment contexts. Table 2 provides a high-level summary of how these goals are addressed in practice. For each privacy property, we list relevant cryptographic primitives, system mechanisms, and example of representative protocols. This mapping serves as a reference point for researchers and practitioners navigating the trade-offs inherent in privacy-preserving digital currency design.
@@
In general, achieving comprehensive privacy requires combining multiple techniques, each with specific trade-offs in terms of confidentiality, unlinkability, and computational demand. For example, blind signatures are lightweight and provide cash-like anonymity but lack unlinkability without centralized coordination. Ring signatures offer anonymity that scales with ring size, but increasing the anonymity set directly impacts computational cost and latency. Encryption, threshold encryption, and commitments facilitate selective auditability but do not provide sender anonymity on their own. ZKPs offer robust privacy guarantees, but they are computationally intensive, may require substantial RAM (e.g., (Almashaqbeh and Solomon, 2022; Liang et al., 2025)), and require careful engineering for practical deployment. One of the well-known ZKP-based systems is zk-SNARK, which makes transactions compact for verification but needs a trusted setup that, if mishandled, could compromise the security of the system. Importantly, as shown in Section 4.4, implementing auditability often requires careful design of ledger data structures and interaction protocols to avoid unintended information leaks while still enabling policy verification. This multidimensional design space underscores the need for a holistic, design-oriented approach: effective privacy-preserving digital currencies must integrate layered technical solutions that balance user privacy with institutional accountability.

### v2

@@ 4.3.6. Plausible Deniability
With plausible deniability, we intend whether a user can deny intentional participation in a transaction, preventing others and adversaries from proving their involvement (e.g., (Almashaqbeh and Solomon, 2022; Ziegeldorf et al., 2018)). Existing works are roughly evenly divided between those supporting this property (e.g., (Noether et al., 2016; Poelstra, 2016; Narula et al., 2018; Kiayias et al., 2022; Tomescu et al., 2022; Wüst et al., 2022; Goodell et al., 2023; Papadoulis et al., 2024)) and those not (e.g., (Maxwell, 2013; Van Saberhagen, 2013; Garman et al., 2017; Bünz et al., 2020; Gross et al., 2021; Baudet et al., 2023; Guan et al., 2022; Yu et al., 2024)). Basically, protocols “on-demand”, such as mixing protocols, subsume an explicit user intention in using the additional service to obtain a privacy feature (such as unlinkability). Likewise, overlay systems implementing an anonymous crypto-asset systems, such as Zerocoin (Miers et al., 2013), Zerocash (Ben Sasson et al., 2014), EZC (Androulaki and Karame, 2014), Zether (Bünz et al., 2020), require explicit intention to convert a native asset to a privacy-preserving one. Conversely, systems that force anonymity or unlinkability as default, readily provide the property of plausible deniability. Among these systems, we mention Monero (Noether et al., 2016), MimbleWimble (Poelstra, 2016), UTT (Tomescu et al., 2022), PEReDi (Kiayias et al., 2022), or Platypus (Wüst et al., 2022), that do not provide the functionality of combining anonymous and public transactions.
@@ 4.4. Auditability
Auditability refers to the ability to examine, track, and verify transactions or balances while preserving a certain degree of privacy. It is a key feature of the third generation of privacy-preserving digital payment systems and, in particular, CBDCs. Two concepts closely related to auditability are: policy and accountability. A policy refers to the rules enforced by the protocol or by users to govern behaviors (e.g., to restrict who can transact and under what conditions). While policies involve ex-ante enforcement, auditability supports post-facto verification. Accountability is a more general concepts related to linking actions to individuals, e.g., through identity-binding mechanisms used by registration authorities to comply with AML/CFT regulations. Full privacy often conflicts with compliance: if transactions are completely opaque, regulators have no oversight (Ballaschk and Paulick, 2021). We identify the following approaches that enable or support auditability: anonymity budget, operating limits, revocable anonymity and user tracing, coin tracing, and verification without disclosure.
@@ 4.4.1. Anonymity Budget
An anonymity budget restricts how many anonymous payments a participant can perform, with each transaction consuming part of it (e.g., (Camenisch et al., 2006; Wüst et al., 2019; Tomescu et al., 2022; ECB, 2019)). Camenisch et al. (Camenisch et al., 2006) propose a bounded-anonymity business model, where a public limit defines how many coins a user may anonymously transfer to a merchant. PRCash (Wüst et al., 2019) instead caps the total anonymous funds a user can receive within a time window. The anonymity budget is represented as a non-spendable output, whose size and duration regulate the flow of anonymous funds. Once exhausted—or if the user chooses non-anonymous transactions—the payee attaches his identity, encrypted under the regulator’s key. UTT (Tomescu et al., 2022) applies the budget to spending, requiring either loss of privacy or auditor intervention once the allocation is exceeded. Auditors can refresh budgets or approve extra transactions after requesting more details. In all these systems, the budget is realized as a special token included in every anonymous yet accountable transaction.
@@ 4.4.2. Operating Limits
Operating limits indicate constraints on transactions and accounts, such as maximum value, user balance, total sent/received funds, or transaction count within a time window. Compliance is generally proven in zero-knowledge. Garman et al. (Garman et al., 2017) extend Zerocash with auditability features, including regulatory closure (i.e., restricting transactions to assets of the same type), spending limits, selective user inquiries, and coin tracing. A central idea is the per-transaction spending limit: transactions above the limit are invalid unless signed by an authority. Another approach uses a per-user counter incremented with each outgoing transaction. The system can then reject transactions beyond a threshold. Gross et al. (Gross et al., 2021) build on this with a two-layer account-based system for implementing a CBDC, where each account tracks epoch turnover (total spent per epoch), proven with ZKPs. Similarly, AQQUA (Papadoulis et al., 2024) maintains multiple counters per account (balance, sent, received), publicly stored as three commitments. A registration authority links users identities to their (possibly, multiple) public keys, while an audit authority check compliance with five policy types: spending/receiving limits, transaction value limits, inactivity over a time period, selective transaction disclosure. During an audit, the authority selects a user by his public key and requests the opening of two historical account state snapshots (for every user’s accounts). After audits, the user randomizes his accounts (and provides a ZKP of valid update), thereby restoring his privacy. Other systems also rely on counters. Xue et al. (Xue et al., 2023) use per-user counters stored in a Merkle tree managed by regulators, which are also in charge of user on-boarding. ZKPs191919Specifically, zk-SNARKs are used for arithmetic statements and Sigma Protocols for algebraic statements. on counters enable policies on total amount and frequency of fund transfers within a time period. When suspicious transactions are detected, the participants’ identities can be recovered by regulators. Chen et al. (Chen et al., 2020) define policies on transaction limits, rates, and value, also verified with ZKPs.
@@ 4.4.3. Revocable Anonymity and User Tracing
User tracing mechanisms enable authorities to discover a user’s identity or link his transactions when required, implementing revocable or conditional anonymity models.202020Revocable anonymity refers to a model where users remain anonymous by default, but their identity can be revealed under certain conditions. User tracing refers to technical mechanisms enabling authorized entities to identify users or link their transactions when required.
@@
Some schemes rely on special digital signatures that expose the sender on double spending (e.g., (Hayes, 1990; Zhang et al., 2019; Beer et al., 2024)). Others embed encrypt identifiers, keys, or metadata into tags, accessible only to tracing authorities (e.g., (Cecchetti et al., 2017; Garman et al., 2017; Zhang et al., 2019; Liu et al., 2022; Lin et al., 2020; Li et al., 2021; Kiayias et al., 2022)). This is often referred to as the trapdoor approach. Many systems also introduce a registration authority that links cryptographic keys to real-world identities (e.g., (Zhang et al., 2019; Liu et al., 2022)). For instance, Cecchetti et al. (Cecchetti et al., 2017) use oblivious RAM to manage transactions across banks, while Liu et al. (Liu et al., 2022) encrypt user identities under the central bank’s key. Zhang et al. (Zhang et al., 2019) rely on a supervision authority with decryption power. Garman et al. (Garman et al., 2017) propose a tracing mechanism where each user receives a unique encryption key: if the user is being traced, the key is a randomized version of the authority’s key; otherwise, it is derived from a null key. Androulaki et al. (Androulaki et al., 2020) encrypt transaction data under the auditors of both sender and receiver, while Xue et al. (Xue et al., 2023) scope tracing by encoding user identities with epoch identifiers. The epoch identifier enables temporal scoping, which limits the impact of key compromising and can enable granular tracing. Other systems derive one-time keys from long-term keys, enabling tracing by recovering the latter. Lin et al. (Lin et al., 2020) let users perform transactions using anonymous keys derived from their public addresses. A manager can use his key and the anonymous key to recover the user public address. Traceable Monero (Li et al., 2021) tags payee one-time keys: an authority can decrypt the tag and perform backward and forward tracing (i.e., identify the accounts that sent/receive funds to/from the current one).
A common concern is the trust concentration in a single tracing authority. To address this, Yu et al. (Yu et al., 2024) split tracing between two authorities via encrypted “marks” requiring joint computation, while PEReDi (Kiayias et al., 2022) (and Sarencheh et al. (Sarencheh et al., 2025)) distribute decryption among multiple parties, requiring quorum approval to identify users and access transaction metadata.
@@ 4.4.4. Coin Tracing
Garman et al. (Garman et al., 2017) enable tracing by marking specific tainted coins, where all derived outputs remain traceable. Each coin carries a fresh encryption key; tracing data for outputs is encrypted under both this key and the authority’s public key, preventing unauthorized tracing by users. To avoid cascade tracing and collusion, exchanges can (verifiably) remove traceability with a public dummy key. Keller et al. (Keller et al., 2021) study de-anonymization in CoinJoin and ring signatures through collaboration with witnesses—i.e., users who confirm their inputs were not part of a mix. As the number of witnesses grows, the anonymity set shrinks, enabling more effective deanonymization.
@@ 4.4.5. Verification without Disclosure
Verification without disclosure lets verifiers confirm policy compliance without accessing private data. These mechanisms rely on cryptographic proofs (usually NIZK) to attest that hidden data satisfies constraints (e.g., (Narula et al., 2018; Wüst et al., 2022; Cecchetti et al., 2017; Bünz et al., 2020; Ma et al., 2021)). The verifier learns only the truth of the statement. Besides the approaches in Operating limits, here we mention those supporting policies beyond basic balance or transfer correctness check (e.g., (Bünz et al., 2020; Miers et al., 2013; Fauzi et al., 2019; Ma et al., 2021; Baudet et al., 2023; Guan et al., 2022; Beer et al., 2024; Papadoulis et al., 2024; Friolo et al., 2024)). zkLedger (Narula et al., 2018) uses Pedersen commitments enabling aggregate computations. zkLedger supports complex auditing via a map-reduce model: banks generate commitments over filtered and transformed data (e.g., checking for non-zero values or squaring for variance) along with NIZKs proofs of correctness, enabling auditors to compute metrics like means and variances. Platypus (Wüst et al., 2022) defines account states with opaque auxiliary data to support custom policy checks. Solidus (Cecchetti et al., 2017) uses PVORM and ZKPs to publicly verify a bank’s transaction log compliance. DSC (Ma et al., 2021) enforces compliance of encrypted transactions with range and authorization proofs.
@@ 4.5. Summary and Takeaways
This section highlights the evolving landscape of privacy goals in digital payment systems. Early approaches focused primarily on anonymity and unlinkability via address obfuscation and mixing, while recent designs emphasize confidentiality and auditability, enabled by advanced cryptographic tools such as ZKPs, homomorphic commitments, and threshold encryption.
Each privacy goal—anonymity, confidentiality, unlinkability, and auditability—relies on distinct mechanisms and entails trade-offs. For instance, unlinkability may hinder auditability, while encrypted transactions that preserve confidentiality can complicate oversight. No single protocol satisfies all privacy and compliance objectives; systems tend to prioritize different goals based on their threat models, trust assumptions, and intended deployment contexts. Table 2 summarizes how privacy goals are addressed in practice, mapping privacy properties to cryptographic primitives, system mechanisms, and representative protocols. This provides a reference for researchers and practitioners navigating the trade-offs in privacy-preserving payment system design.
@@
Privacy-preserving payment systems typically combine multiple cryptographic tools, each with distinct trade-offs. Blind signatures are lightweight but centralized; ring signatures provide scalable anonymity at latency cost; threshold encryption and commitments support auditability but not anonymity; and ZKPs (e.g., zkSNARKs, zkSTARKs) offer strong guarantees, but they are computationally intensive, may require substantial RAM (e.g., (Almashaqbeh and Solomon, 2022; Liang et al., 2025)), and require careful engineering for practical deployment. One of the well-known ZKP-based systems is zk-SNARK, which makes transactions compact for verification but needs a trusted setup that, if mishandled, could compromise the security of the system. Importantly, as shown in Section 4.4, implementing auditability often requires careful design of ledger data structures and interaction protocols to avoid unintended information leaks while still enabling policy verification. This multidimensional design space underscores the need for a holistic, design-oriented approach: effective privacy-preserving digital payment systems must integrate layered technical solutions that balance user privacy with institutional accountability.